OpenStudy (anonymous):

please help me with this question:prove that 7log16\15+5log25\24+3 log81\80=log2

OpenStudy (anonymous):

what is the 9th term if the nth term is n(n+1)\3 )

OpenStudy (anonymous):

what is the arithmetic mean of 34 and 44?

OpenStudy (michele_laino):

Question #1 hint: \[\begin{gathered} \log \left( {\frac{{16}}{{25}}} \right) = \log 16 - \log 25 = \log \left( {{2^4}} \right) - \log \left( {{5^2}} \right) = 4\log 2 - 2\log 5 \hfill \\ \log \left( {\frac{{25}}{{24}}} \right) = \log 25 - \log 24 = \log \left( {{5^2}} \right) - \log \left( {{2^3} \cdot 3} \right) = 2\log 5 - 3\log 2 - \log 3 \hfill \\ \log \left( {\frac{{81}}{{80}}} \right) = \log 81 - \log 80 = \log \left( {{3^4}} \right) - \log \left( {{2^4} \cdot 5} \right) = 4\log 3 - 4\log 2 - \log 5 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

question #3 what is: \[\frac{{34 + 44}}{2} = ...?\]

OpenStudy (anonymous):

oho its not 16/25 its 16/15

OpenStudy (michele_laino):

oops.. you are right! \[\begin{gathered} \log \left( {\frac{{16}}{{15}}} \right) = \log 16 - \log 15 = \log \left( {{2^4}} \right) - \log \left( {3 \cdot 5} \right) = 4\log 2 - \log 3 - \log 5 \hfill \\ \log \left( {\frac{{25}}{{24}}} \right) = \log 25 - \log 24 = \log \left( {{5^2}} \right) - \log \left( {{2^3} \cdot 3} \right) = 2\log 5 - 3\log 2 - \log 3 \hfill \\ \log \left( {\frac{{81}}{{80}}} \right) = \log 81 - \log 80 = \log \left( {{3^4}} \right) - \log \left( {{2^4} \cdot 5} \right) = 4\log 3 - 4\log 2 - \log 5 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

thank you so much

OpenStudy (michele_laino):

thank you!

OpenStudy (anonymous):

I've another doubt please clarify !what is A.G.P(arithmetic geometric progression).

OpenStudy (michele_laino):

keep in mind that there are two progressions, namely 1) the arithmetic progression 2) the geometric progression

OpenStudy (michele_laino):

we have an arithmetic progression when the difference between one term and its consecutive term (previous or subsequent) is constant

OpenStudy (michele_laino):

whereas we hae a geometri progression, where the quotient between one term and its consective term (previous or subsequent) is costant

OpenStudy (anonymous):

thank you for explaining me about progression now I can manage progressions